Paul Coddington

Northeast Parallel Architectures Center 111 College Place, 3-217 CST Syracuse University

July 2, 1997

Professor Geoffrey Fox Northeast Parallel Architectures Center 111 College Place, 3-217 CST Syracuse University

Dear Professor Fox,

This is to confirm that I will be resigning my position with the Northeast Parallel Architectures Center at Syracuse University effective 31 July 1997, in order to take up a new position in Australia.

My forwarding address will be:

Paul Coddington

Department of Computer Science The University of Adelaide Adelaide, SA 5005 Australia

Email: paulc@cs.adelaide.edu.au

Thank you very much for your support during my time at Syracuse.

Yours sincerely,

Paul Coddington